OpenStudy (anonymous):

you have a part time job at a local restraunt. One evening you work 3 hours and earn 20.25. on Saturday you work 6 hours and earn 40.50. determine your hourly wage, the rate of the change.

OpenStudy (roadjester):

Is this a calc problem? it mentions "rate of change" which is a derivative in calculus

OpenStudy (anonymous):

rate of change just means compute the slope of \[(3,20.25), (6,40.5)\]

OpenStudy (roadjester):

although, i would just say, if it's a simple problem, make it (amount earned)/(number of hours worked)

OpenStudy (roadjester):

so you get dollars per hour which is still a rate of change

OpenStudy (anonymous):

a really dumb question since unless you get paid for just showing up you might as well divide \[\frac{20.25}{3}\]

OpenStudy (roadjester):

6.75 dollars/hour

OpenStudy (anonymous):

no reason to compute a slope for sure. unless you get money for clocking in. you make $6.75 per hour

OpenStudy (roadjester):

as long as you have something/something, it's a rate of change

OpenStudy (amistre64):

20.25 ------- ; it the second data point wasnt a scalar of the first then wed have to 3 do a little more work to find the average hourly pay

OpenStudy (anonymous):

It's about 6.80

OpenStudy (amistre64):

if I work for 3 hours and make 20.25; then I expect to work for 6 hours at the same pay scale and make 40.50 since I worked twice as long

OpenStudy (amistre64):

we can put in a 0 pay for 0 hours work and see if it matches up to a line or a quad :)

OpenStudy (amistre64):

pay = 6.75h

OpenStudy (anonymous):

this is a dumb question yes

OpenStudy (amistre64):

(6,40.50) (3,20.25) --------- (3,20.25); slope = y/x = 20.25/3 = 6.75

OpenStudy (anonymous):

wage = pay / time worked. they gave us

OpenStudy (amistre64):

not really 'dumb'. Its a simple question yes; but not a 'dumb' question perse..

OpenStudy (anonymous):

sorry guys, im not a math person....